OpenStudy (anonymous):

evalute the function as indicated and simplify f(x)=3-7x (c)f(t)+f(-2)

OpenStudy (turingtest):

ummm..... (c)f(t)=3c-7ct f(-2)=3+14=17 + -------------- so 17+3c-7ct if you wrote it right, that seems like a weird question to me.

OpenStudy (anonymous):

no c is not part of the question

OpenStudy (turingtest):

so f(t)+f(-2)=? is the question

OpenStudy (anonymous):

yeah can you show me how to do it?

OpenStudy (turingtest):

yeah f(t)=3-7t f(-2)=3-7(-2)=3+14=17 f(t)+f(-2)=3-7t+17=-7t+20 helpful?

OpenStudy (anonymous):

yeah thanks